OpenStudy (hitaro9):

Determine if 83 divides 2^41 +1 using quadratic results of quadratic residues

OpenStudy (hitaro9):

I can pretty easily go through and set up a congruent like x = 2^41+1 mod 83 and reduce that to x = 0 to show that there's no remainder But with quadratic residues I'm not sure what residues to point out. If I set up x-1 = 2^41 mod 83 I find that this mirrors the euler's criterion ie (2/41) = 2^41 mod 83 But I'm not sure of the relevance.

ganeshie8 (ganeshie8):

Little fermat gives us \[2^{82}\equiv1 \pmod{83}\] or \[2^{82}-1\equiv 0 \pmod{83}\] factor and get \[(2^{41}-1)(2^{41}+1)\equiv 0\pmod{83}\]

ganeshie8 (ganeshie8):

Because \(83\) is a prime, it follows that \(2^{41}-1\equiv 0\pmod{83}\) or \(2^{41}+1\equiv 0\pmod{83}\)

ganeshie8 (ganeshie8):

Notice that first congruence is same as the euler's criterion if the first congruence held, then \(2\) is a quadratic residue of \(83\)

ganeshie8 (ganeshie8):

if we could show that \(2\) is "not" a quadratic residue of \(83\), then thats same as proving that the first congruence does not hold. that lets us to conclude that the second congruence holds.

OpenStudy (hitaro9):

Oohhh okay

OpenStudy (hitaro9):

That makes a lot of sense

ganeshie8 (ganeshie8):

so try and show that "2" is not a quadratic residue of 83

OpenStudy (hitaro9):

Right

OpenStudy (hitaro9):

I'm gunna go through and do that since I get it from there. I've got 1 last problem I was stuck on after this. I know I've bothered you a whole bunch today, but would you mind looking at it?

ganeshie8 (ganeshie8):

sure, but wiat do u know how to test whether "2" is a quadratic residue or not ?

ganeshie8 (ganeshie8):

You may use below to evaluate the legendre symbol (2/p) :\[(2/p) = (-1)^{(p^2-1)/8}\]

OpenStudy (hitaro9):

There was a book problem asking a similar problem, I was just going to take a look at that

OpenStudy (hitaro9):

Right that looks familiar.

ganeshie8 (ganeshie8):

Okay

OpenStudy (hitaro9):

@ganeshie8 Sorry, lost connection to openstudy :c

OpenStudy (hitaro9):

The question is: Let p>= 5 havequadratic residues a1 a2.... a(p-1)/2 prove that p divides the sum from i = 1 to (p-1)/2 of ai

OpenStudy (hitaro9):



OpenStudy (hitaro9):

^There's a terrible drawing if my written way of putting it wasn't clear

ganeshie8 (ganeshie8):

Notice that the quadratic residues of \(p\) are precicely the integers from set:\[\{1^2, 2^2, ...,((p-1)/2)^2\}\] Therefore the sum \[\sum\limits_{i=1}^{(p-1)/2}a_i \] is same as \[\sum\limits_{i=1}^{(p-1)/2}i^2 \]

ganeshie8 (ganeshie8):

whats the sum of squares of first (p-1)/2 integers ?

OpenStudy (hitaro9):

I'm sorry, I'm just not seeing it

OpenStudy (hitaro9):

Oh wait. There's this sum thing. One moment

OpenStudy (hitaro9):

Yeah, the sum from 1 to n of i^2 is n(n+1)(2n+1)/6

OpenStudy (hitaro9):

Is that what you're hinting at? Cause I imagine that simplifies out nicely.

ganeshie8 (ganeshie8):

yes, plugin n=(p-1)/2

OpenStudy (hitaro9):

Alright then, that's pretty clever. Thank you.

OpenStudy (hitaro9):

You're always super helpful, I really do appreciate it.

ganeshie8 (ganeshie8):

np:)